OpenStudy (anonymous):

Can someone double check my work? Simplify the following expression: (5x2 + 3x + 4) − (2x2 − 6x + 3). If the final answer is written in the form Ax2 + Bx + C, what is the value of B? I got 3x^2 - 3x + 7, so Bx = -3x??

jimthompson5910 (jim_thompson5910):

When you distribute, you need to distribute that negative to EVERY term inside the second parenthesis.

jimthompson5910 (jim_thompson5910):

So (5x^2 + 3x + 4) − (2x^2 − 6x + 3) becomes 5x^2 + 3x + 4 - 2x^2 + 6x - 3 Notice how it's +6x and NOT -6x

OpenStudy (anonymous):

so combine the terms now?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

is the answer 9x?

jimthompson5910 (jim_thompson5910):

5x^2 - 2x^2 = 3x^2 3x + 6x = 9x 4 - 3 = 1 So the simplified expression is 3x^2 + 9x + 1 It's in the form Ax^2 + Bx + C So we can see that Bx = 9x, which means B = 9

OpenStudy (anonymous):

thank you! (:

jimthompson5910 (jim_thompson5910):

you're welcome